OpenStudy (anonymous):

Solve: -2x - 15 = 6x + 9

OpenStudy (ja1):

Ok, well we want to isolate the variable. Do you know how to start off?

OpenStudy (anonymous):

X=-3 -2x-15=6x+9 Since 6x contains the variable to solve for, move it to the left-hand side of the equation by subtracting 6x from both sides. -2x-15-6x=9 Since -2x and -6x are like terms, subtract 6x from -2x to get -8x. -8x-15=9 Since -15 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 15 to both sides. -8x=15+9 Add 9 to 15 to get 24. -8x=24 Divide each term in the equation by -8. -(8x)/(-8)=(24)/(-8) Simplify the left-hand side of the equation by canceling the common factors. x=(24)/(-8) Simplify the right-hand side of the equation by simplifying each term. x=-3